Part 1. Random Events 1. A square has a side of 21 cm and has a circle with radius equal to 7 cm within it. Find the probability of a dart hitting the circle if it hits the square.
Answer #1

7CM 21 CMs

The are of Circle:

A = π * r2ー3.14 * 72-153.86m (Here Pi=3.14)

The area of Square:

A = r2-212-44177,-

The probability of hitting the circle: The area of Circle/ The area of square= 153.86/441= 0.348889


The probability of P= 0.348889